OpenStudy (anonymous):

The length of a rectangle is 3 units and its width is square root of 3 unit. Is the area of the rectangle rational or irrational

OpenStudy (anonymous):

@Catseyeglint911 @superhelp101

OpenStudy (anonymous):

I know the answer is irrational, I just don't know how to get the actual answer

OpenStudy (anonymous):

length times width = the area of a rectangle.

OpenStudy (anonymous):

So 3 times the square root of 3 = the area.

OpenStudy (anonymous):

okay but how would i multiply the radical

OpenStudy (anonymous):

Do you have a calculator on you?

OpenStudy (anonymous):

I think it would be \[3*3\frac{ 1 }{ 2 }\]

OpenStudy (anonymous):

yes I o have a calculator

OpenStudy (anonymous):

square root of 3 = 1.73205080757 3*1/2 is half of 3 not the square root. Square root: a number that produces a specified quantity when multiplied by itself. For example 3 * 2 = 6 but 3 * 3 = 9 which is why 3 is the square root of 9. Make sense?

OpenStudy (acxbox22):

3 is length sqrt 3 is width a=lw \[A=3\sqrt{3}\]

OpenStudy (anonymous):

So the area would be 9?

OpenStudy (acxbox22):

no, the area would be \[3\sqrt{3}~~units^2\]

OpenStudy (anonymous):

I got 5.2

OpenStudy (acxbox22):

if the area was 9 then the are would be rational which it isnt

OpenStudy (acxbox22):

5.2 is the rounded answer but 3sqrt3 is the exact answer

OpenStudy (anonymous):

acxbox22 is right. and by the way my answer is rounded.

OpenStudy (anonymous):

ohhhh....... that makes sense now

OpenStudy (anonymous):

@Catseyeglint911 how does \[3\sqrt{3}\] make 5.2

OpenStudy (acxbox22):

round the sqrt of 3 and multiply it by 3

OpenStudy (anonymous):

Just multiply 3 by the square root of 3

OpenStudy (anonymous):

oops typed too slow. XD

OpenStudy (anonymous):

ohhhh... THAT MAKES A LOT O SENSE XD